# Human-Aware 3D Scene Generation with Spatially-constrained Diffusion Models

### v1

@@ Introduction
Baselines. We compare our method with MIME (Yi et al. 2023) and DiffuScene (Tang et al. 2023) using their official implementations. MIME is a transformer-based autoregressive method for human-aware scene generation, while DiffuScene is a diffusion-based method that learns 3D scene distributions without conditioning on the floor plan. For a more fair comparison, we adapt DiffuScene on the 2D-floor plan with the free-space mask to enhance its perception of human motions. All methods are trained on our calibrated 3D FRONT HUMAN dataset.
@@
Evaluation Metrics. Following (Yi et al. 2023), we evaluate all methods on the plausibility of human-object interaction and the realism of the generated scenes. As suggested in (Yi et al. 2023), we use the 3D IoU score to measure the plausibility of human-object interaction, by calculating the intersection ratio between input contact bounding boxes and generated objects. Additionally, we employ $\mathbf{Col_{mot}}$ (Yi et al. 2023) to evaluate collision between generated objects and free-space humans, and $\mathbf{Col_{obj}}$ (Yang et al. 2024) to represent the collision rate between objects in the generated scene. To measure the violation of the room layout, we compute the collision rate between generated objects and the areas outside the floor plan, denoted as $\mathbf{R_{out}}$. Furthermore, following (Paschalidou et al. 2021), we calculate the Fréchet inception distance (FID) score and the category KL divergence (CKL) between generated scenes and real scenes to measure the realism of generated scenes. All these evaluation experiments are conducted on the test split of the calibrated 3D FRONT HUMAN dataset.
Figure 5: Ablation on spatial collision guidance functions. The top row shows the scenes generated without guidance (w/o), with red boxes indicating constraint violations, while the bottom row shows scenes generated with guidance (w), where green boxes highlight the improvements achieved by the proposed guidance.
@@
0.082 0.003 0.020
0.034 0.008 0.023
0.100 0.002 0.021
0.084 0.004 0.008
0.051 0.006 0.016
Figure 6: Ablation study on varying numbers of free-space and contact humans. (a) As the number of free-space humans increases, SHADE generates fewer objects in the scene. (b) Providing more contact humans as input, SHADE generates more occupied objects to interact with them.
@@
Figure 4 exhibits the generation ability of our method and baselines for different room types. Since DiffuScene only considers floor plan and free-space information, it fails to generate appropriate objects to interact with contact humans. Additionally, DiffuScene also generates objects colliding with free-space humans. Both MIME and our method can generate reasonable objects to support various contact humans, such as placing a bed under a sitting human or a sofa under a lying human. However, MIME still generates objects in free-space humans or outside the floor plan (see the second column of Fig. 4). In contrast, our method generates more plausible scenes that can avoid colliding with free-space humans and room boundaries, as well as prevent object overlap.
These observations are validated by the quantitative comparisons of various evaluation metrics presented in Table 2. Our method outperforms MIME in 3D IoU scores for both the bedroom and living room and performs comparably in the dining room. Furthermore, our method significantly reduces motion collisions ($\mathbf{Col_{mot}}$) and object collisions ($\mathbf{Col_{obj}}$) across all room types when compared to the baselines. Although MIME achieves better FID scores, our method consistently shows better CKL scores, indicating better alignment with the categorical distribution of real data. Overall, these results demonstrate the effectiveness of our framework.
@@
Following MIME (Yi et al. 2023), we test SHADE on a real dataset of human motion to evaluate the generalization of our method. We use 5 input motions from the PROXD (Hassan et al. 2019) dataset and generate 10 scenes for each motion sequence. The quantitative results of our method and the baselines are reported in Table 3. Note that all methods are not fine-tuned on the PROXD dataset. For human-object interaction plausibility, our method achieves a 3D IoU score of 0.881, outperforming DiffuScene and being comparable to MIME. In terms of scene plausibility, our method surpasses the other methods on the motion collision metric $\mathbf{Col_{mot}}$ and the object collision metric $\mathbf{Col_{obj}}$, indicating that our method can generate more realistic scenes with fewer human-object and object-object collisions. Figure 1 presents the qualitative visualization of all models’ generations.
Ablation Studies
@@
Inference Guidance. We investigate the impact of each spatial collision guidance function on the bedroom and present the results in Tab. 4. Compared to scene generation without any guidance function, incorporating the motion collision avoidance function reduces the $\mathbf{Col_{mot}}$ metric to 0.034, demonstrating its effectiveness. Similar conclusions can be drawn from the results of the other two spatial collision guidance functions. It is noteworthy that these collision-based guidance functions can negatively affect each other. For example, while the room boundary constraint function improves the $\mathbf{R_{out}}$ metric, it can degrade the other metrics. This is reasonable because room boundary guidance pushes objects within the floor plan, potentially leading lead to increased collisions with free-space humans ($\mathbf{Col_{mot}}$) and other objects ($\mathbf{Col_{obj}}$). To balance these effects, we integrate all spatial collision functions, thereby achieving better overall performance in scene plausibility. Fig. 5 provides a qualitative visualization of the effect of each spatial collision guidance function. The improvements shown in the second row of Fig. 5 confirm that our collision-based guidance functions can significantly enhance the plausibility of 3D scenes.
Number of Input Humans. In Fig. 6, we further investigate the impact of input humans on scene generation by varying the number of free-space humans and contact humans provided as input. Our qualitative results indicate that as the density of free-space humans increases, our method generates fewer objects in the scenes. Additionally, when given more contact humans, we produce more occupied objects to interact with them. These findings demonstrate the flexible scene generation capabilities of our method, influenced by the varying numbers of input humans.

### v2

@@ Introduction
Baselines. We compare our method with MIME (Yi et al. 2023) and DiffuScene (Tang et al. 2023) using their official implementations. MIME is an autoregressive method for human-aware scene generation, while DiffuScene is a diffusion-based method that learns 3D scene distributions without conditioning on the floor plan. For a more fair comparison, we adapt DiffuScene on the 2D-floor plan with the free-space mask to enhance its perception of human motions. All methods are trained on our calibrated datasets.
@@
Evaluation Metrics. Following (Yi et al. 2023), we evaluate all methods on the plausibility of human-object interaction and scene realism. To measure the plausibility of human-object interaction, we estimate the 3D IoU score by calculating the intersection ratio between input contact bounding boxes and generated objects. We also employ $\mathbf{Col_{mot}}$ (Yi et al. 2023) to evaluate collision between generated objects and free-space humans, and $\mathbf{Col_{obj}}$ (Yang et al. 2024) to represent the collision rate between objects and objects. To measure the violation of the room layout, we compute the collision rate between generated objects and the areas outside the floor plan, denoted as $\mathbf{R_{out}}$. For realism, we calculate the Fréchet inception distance (FID) score and the category KL divergence (CKL) (Paschalidou et al. 2021) between generated and real scenes. All evaluation experiments are conducted on the test split of the calibrated datasets.
Network Architecture. As depicted in Fig. 2, we implement SHADE with a transformer-based architecture composed of four DiT blocks with adaLN-Zero (Peebles and Xie 2022). As input, all object parameters are encoded by a fully connected network, while the time step $t$ is encoded using a positional embedding (Tancik et al. 2020). To make the scene generation aware of human motion sequences and room layout, contact humans, the floor plan $\mathcal{F}$ and the free-space mask $\mathcal{FS}$ are encoded by a fully connected network and PointNet (Qi et al. 2017), respectively. These embeddings adjust the scale and shift parameters of each adaLN-Zero block. See more details in Sup. Mat.
Implementations. We train our human-aware scene diffusion model with the default settings in (Ho, Jain, and Abbeel 2020b), and reduce the number of diffusion steps to $T$=200 from 1000 for speeding up the sampling process. See more details in Sup. Mat. At inference time, we use the DDPM sampler (Ho, Jain, and Abbeel 2020b) to obtain the object properties. Subsequently, we retrieve the closest CAD model from the 3D-FUTURE dataset (Fu et al. 2021) based on the class label and bounding box size of the generated objects.
0.043 0.002 0.013
0.034 0.002 0.013
0.045 0.001 0.014
0.036 0.003 0.001
0.027 0.003 0.001
8
@@
Figure 4 exhibits the generation ability of our method and baselines for different room types. Since DiffuScene only considers floor plan and free-space information, it fails to generate appropriate objects to interact with contact humans. Additionally, DiffuScene generates objects colliding with free-space humans. Both MIME and our method can generate reasonable objects to support various contact humans, such as placing a bed under a sitting human or a sofa under a lying human. However, MIME still generates objects in free-space humans or outside the floor plan (see the second column of Fig. 4). In contrast, our method generates more plausible scenes that can avoid colliding with free-space humans and respect room boundaries, as well as prevent object overlap. These observations are validated by the quantitative comparisons presented in Table 2. Our method outperforms the baselines in 3D IoU scores and significantly reduces motion collisions ($\mathbf{Col_{mot}}$), object collisions ($\mathbf{Col_{obj}}$) and room layout violations ($\mathbf{R_{out}}$) across all room types. Additionally, our method achieves the best FID scores in the Bedroom and Living room settings, and the best CKL scores in the Bedroom and Dining room settings, indicating that our method generates more realistic scenes. Overall, these results demonstrate the effectiveness of SHADE.
@@
Following MIME (Yi et al. 2023), we test SHADE on a real dataset of human motion to evaluate the generalization of our method. We use 5 input motions from the PROXD (Hassan et al. 2019) dataset and generate 10 scenes for each motion sequence. The quantitative results of our method and the baselines are reported in Table 3. All methods are not fine-tuned on the PROXD dataset. For human-object interaction plausibility, our method achieves a 3D IoU score of 0.881, outperforming MIME and DiffuScene. In terms of scene plausibility, our method surpasses the baseline methods on the motion collision metric $\mathbf{Col_{mot}}$ and the object collision metric $\mathbf{Col_{obj}}$, indicating that our method generates more realistic scenes with fewer human-object and object-object collisions. Figure 1 presents the qualitative visualization of all models’ generations.
Figure 5: Ablation on spatial collision guidance functions. The left column shows scenes generated without guidance, with red boxes indicating constraint violations. The right column shows scenes with guidance, where green boxes highlight the improvements.
Ablation Study on Inference Guidance
@@
We investigate the impact of each spatial collision guidance function on the bedroom and present the results in Tab. 4. Compared to scene generation without any guidance function, incorporating the motion collision avoidance function reduces the $\mathbf{Col_{mot}}$ metric to 0.034 from 0.043, demonstrating its effectiveness. Similar conclusions can be drawn from the results of the other two spatial collision guidance functions. It is noteworthy that these collision-based guidance functions can negatively affect each other. For example, while the room boundary constraint function improves the $\mathbf{R_{out}}$ metric, it can degrade the other metrics. This is reasonable because room boundary guidance pushes objects within the floor plan, potentially leading to increased collisions with free-space humans ($\mathbf{Col_{mot}}$) and other objects ($\mathbf{Col_{obj}}$). To balance these effects, we integrate all spatial collision functions, thereby achieving better overall performance in scene plausibility. Fig. 5 provides a qualitative visualization of the effect of each spatial collision guidance function. The improvements shown in Fig. 5 confirm that our collision-based guidance functions can significantly enhance the plausibility of 3D scenes.